# matrix,arrows,decorations.pathmorphing

### v2

@@ 2.5. Four lemmas
Let $1\to H\to G\to L\to 1$ be a short exact sequence of groups and let $f$ be an automorphism of $H$ such that:
@@
For an element $f\in\operatorname{Aut}(H)$, we denote by $\bar{f}$ its image in $\operatorname{Out}(H)$. Let then $\mathrm{Comp}(L,H)$ be the closed subgroup of $\operatorname{Aut}(L)\times\operatorname{Aut}(H)$ formed by the pairs $(\psi,f)$ such that, for all $\alpha\in L$, there holds (in $\operatorname{Out}(H)$):
@@
Since, by hypothesis (i), the center of $H$ is trivial, according to Wells’ exact sequence (cf. Theorem in [25]), there is a canonical isomorphism:
@@
2.6. Connectedness of various curve complexes
Before to proceed with the proof of Proposition 2.1, we need to prove that some curve complexes are connected.
Let $C_{b}(S_{0,n})$ be the curve complex defined as the full subcomplex of the complex of curves $C(S_{0,n})$, for $n\geq 4$, whose vertices consist of isotopy classes of simple closed curves on $S_{0,n}$ which bound a $2$-punctured disc:
Lemma 2.12.
For $n\geq 5$, the simplicial complex $C_{b}(S_{0,n})$ is connected.
Proof.
This can be proved by the same argument which proves the connectivity of the standard curve complex (cf., for instance, the proof of Theorem 4.3 in [9]). We use induction on the geometric intersection number of two simple closed curves $a$ and $b$ which bound a $2$-punctured disc on $S_{0,n}$. When they are disjoint there is nothing to prove. Let us then assume that $a$ and $b$ have geometric intersection $i(a,b)>0$. We claim that there is a simple closed curve $c$ (bounding a $2$-punctured disc) such that $i(a,c)=0$ and $i(c,b)<i(a,b)$.
Now, $c^{\prime}$ bounds a $k$-punctured disk $D$, where $k\geq 2$, which does not contain $a$. If $k=2$, then we take $c=c^{\prime}$ and we are done. Otherwise, we take for $c$ the boundary of a $2$-punctured disc $D^{\prime}$ contained in $D$ such that $c$ is the union of an arc contained in $c^{\prime}$ and an arc which is either part of $b\cap D$ or is disjoint from $b$. In both cases, we have that $i(c,b)\leq i(c^{\prime},b)<i(a,b)$ and, obviously, $i(a,c)=0$. ∎
Let then $C_{0b}(S_{1,n})$ be the full subcomplex of the curve complex $C(S_{1,n})$ whose vertices are isotopy classes of either nonseparating simple closed curves or simple closed curves which bound a $2$-punctured disc on $S_{0,n}$:
Lemma 2.13.
For $n\geq 2$, the simplicial complex $C_{0b}(S_{1,n})$ is connected.
Proof.
We proceed as in the genus zero case. Here, the curve $c$ is either nonseparating or it bounds a $2$-punctured disk but the proof above works without any essential change. ∎
2.7. Proof of Proposition 2.1 for $g(S)=0$
We proceed by induction on $n\geq 5$. The case $n=5$ was proved above and serves as base for the induction. Let us then assume that Proposition 2.1 holds for $S_{0,n-1}$ and let us prove it for $S_{0,n}$.
Lemma 2.14.
Proof.
By Theorem 6.7 in [6], for all hyperbolic surfaces $S$, there is a natural monomorphism:
which is induced by the identification of the vertices of ${\check{C}}_{P}(S)$ with the facets of ${\check{C}}(S)$. For $d(S)>1$, one further observes that the continuous action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on the profinite set of Farey subgraphs of ${\check{C}}_{P}(S)$, which are parameterized by the $(d(S)-2)$-simplices of ${\check{C}}(S)$, does indeed induce a continuous action on this profinite set of $(d(S)-2)$-simplices, which is compatible with the action on the facets of ${\check{C}}(S)$ considered above. One then shows that an automorphism of ${\check{C}}(S)$ can be reconstructed from these data.
Thus, after possibly composing the given automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))$, we can assume that its image $\check{\Theta}_{P}(\phi)$ in $\operatorname{Aut}({\check{C}}(S_{0,n}))$ preserves the 00-simplex $\{{\gamma}\}\in{\check{C}}(S_{0,n})$ and so $\phi$ preserves the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{0,n})$. Since, by Lemma 2.8, we have that $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, from the induction hypothesis, it follows that, if the automorphism $\phi$ preserves the orientation of some profinite Farey subgraph of $L_{\gamma}$, then $\phi$ preserves the orientation of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$. ∎
By Lemma 2.12, the curve complex $C_{b}(S_{0,n})$ is connected. Thus, there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S_{0,n}$ bounding a $2$-punctured disc such that any two representatives ${\sigma}$ and ${\sigma}^{\prime}$ of the set of orbits $O(S_{0,n})$ are contained in a chain $\operatorname{Lk}({\gamma}_{1}),\dots,\operatorname{Lk}({\gamma}_{k})$ with the property that the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $11\leq i\leq k-$ 1, contains at least a profinite Farey subgraph. Lemma 2.14 and a simple induction then imply that an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$, which preserves the orientation of ${\widehat{\operatorname{F}}}_{\sigma}$, also preserves the orientation of ${\widehat{\operatorname{F}}}_{{\sigma}^{\prime}}$. This completes the proof of Proposition 2.1 for $g(S)=0$.
2.8. Proof of Proposition 2.1 for $g(S)\geq 1$
Let us first consider the case $g(S)=1$. Here, we need to use the curve complex $C_{0b}(S_{1,n})$ instead of the curve complex $C_{b}(S_{0,n})$ and Lemma 2.13 instead of Lemma 2.12. We then proceed by induction on $n\geq 2$. The base of the induction is provided by the case $S=S_{1,2}$ proved above. The rest of the argument proceeds similarly to the case $g(S)=0$ where, besides the induction hypothesis, we also use the case $g(S)=0$ of Proposition 2.1 proved in Section 2.7.
For $g(S)\geq 2$, we proceed by induction on the genus where the base of the induction is the genus $1$ case proved above. The relevant curve complex here is the complex of nonseparating curves $C_{0}(S)$, which, for $g(S)\geq 2$, (cf. Theorem 4.4 in [9]) is connected. The rest of the argument proceeds as in the previous cases.
2.9. A rigidity criterion
From Proposition 8.2 in [6] and Proposition 2.1, it follows that, for $d(S)>1$ and $S\neq S_{1,2}$, there is a natural isomorphism:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S)).$
For $S=S_{1,2}$, the group $\operatorname{Aut}({\check{C}}_{P}(S_{1,2}))$ must be replaced with the subgroup of those automorphisms preserving the set of nonseparating curves. We will show that this isomorphism implies a characterization of those elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ which are induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$. Before we state the result, we need to make a remark.
In the group-theoretic realization of the procongruence curve complex ${\check{C}}(S)$ which we described in Section 2.5, the vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ are identified with the set $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-1}}$ of inertia subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ of maximal rank. The natural faithful continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the curve complex ${\check{C}}(S)$ (cf. Lemma 2.9) then induces a continuous faithful action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the vertex set ${\check{C}}_{P}(S)_{0}$ of the procongruence pants complex. We then have:
Theorem 2.15.
@@ 2.9. A rigidity criterion
Let $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ be an element satisfying the hypotheses of the theorem. Since, for $S=S_{0,5}$, the action of ${\check{\Gamma}}(S)$ on the oriented edges of ${\check{C}}_{P}(S)$ is transitive, after composing with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S))$, we can assume that $f$ fixes the vertices of the edge $\{v_{0},v_{1}\}$ of ${\check{C}}_{P}(S)$. For the same reason, for any edge $\{\alpha_{0},\alpha_{1}\}$ of ${\check{C}}_{P}(S)$, there is an element $x\in{\check{\Gamma}}(S)$ such that $\alpha_{i}=x\cdot v_{i}\cdot x^{-1}$, for $i=0,1$.
Let us consider the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to{\check{\Gamma}}(S)\to\Sigma_{n}\to 1$, where we put $n:=n(S)$. From Corollary C in [11], it follows that the outer representation associated to this short exact sequence identifies $\Sigma_{n}$ with a normal subgroup of $\operatorname{Out}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$. Since $\mathrm{P}{\check{\Gamma}}(S)$ is center free, from Lemma 2.11, it follows that the given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ extends to an automorphism of ${\check{\Gamma}}(S)$, which we also denote by $f$, so that we have:
$f(\alpha_{i})=f(x\cdot v_{i}\cdot x^{-1})=f(x)\cdot v_{i}\cdot f(x)^{-1},\hskip 14.22636pt\mbox{ for }i=0,1,$
and then $\{f(\alpha_{0}),f(\alpha_{1})\}=\operatorname{inn}f(x)(\{v_{0},v_{1}\})\in{\check{C}}_{P}(S)$. Therefore, the continuous action of the automorphism $f$ on the profinite set of vertices of ${\check{C}}_{P}(S)$ extends to a continuous action on the procongruence pants complex ${\check{C}}_{P}(S)$. The conclusion then follows from Proposition 2.1.
@@
To deal with this case, we need first to prove the following lemmas:
@@
With the notation of Section 2.5, for ${\gamma}$ a simple closed curve on $S_{1,2}$, by Corollary 4.12 in [6], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}$ and, by Theorem 4.10 in [6], there are exact sequences:
$1\rightarrow\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}\to\{\pm 1\}\hskip 14.22636pt\mbox{ and }\hskip 14.22636pt1\rightarrow\hat{\mathrm{I}}_{\gamma}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\to 1$
@@
After possibly composing with an inner automorphism, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ is such that ${\gamma}^{\prime}:=f({\gamma})$ also belongs to $C(S_{1,2})_{0}\subset{\check{C}}(S)_{0}$. Since $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$ identifies with the subgroup of the centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})$ topologically generated by profinite Dehn twists, we have that $f(\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}})=\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}^{\prime}}$. By Theorem 4.14 in [6], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$, so that $f$ induces an isomorphism $\bar{f}\colon\thinspace\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\mathrm{P}{\check{\Gamma}}((S_{1,2})_{{\gamma}^{\prime}})$.
@@
For ${\gamma}$ separating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\cong\widehat{\operatorname{SL}(2,\mathbb{Z})}$ while, for ${\gamma}$ nonseparating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})$ is a free group in two generators. The latter profinite group is torsion free while the former is not. Thus, ${\gamma}^{\prime}$ has the same topological type of ${\gamma}$.
@@
This proves the first part of the lemma. Let us then observe that, by Lemma 2.5, $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ identifies with the normal subgroup of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ topologically generated by squares of nonseparating Dehn twists. By the previous part of the proof, elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ preserve this subgroup and so there is a homomorphism as claimed in the lemma. The fact that this is injective follows from the fact that the monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{1,2}))$ (cf. Lemma 2.9) factors through it and the monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{0,5}))$, via the isomorphism ${\check{C}}(S_{1,2})\cong{\check{C}}(S_{0,5})$. ∎
@@
Since all the groups involved are center free, for $n\geq 4$, there are a series of natural isomorphisms:
@@
$\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\right.\cong{\check{\Gamma}}(S_{0,n})\left/\mathrm{P}{\check{\Gamma}}(S_{0,n})\right.\cong\Sigma_{n}.$

### v3

@@ 2.6. Connectedness of various curve complexes
  1. (i)
    For $n\geq 5$, the simplicial complex $C_{b}(S_{0,n})$ is connected.
  2. (ii)
    For $n\geq 2$, the simplicial complex $C_{0b}(S_{1,n})$ is connected.
Proof.
(i): This can be proved by the same argument which proves the connectivity of the standard curve complex (cf., for instance, the proof of [13, Theorem 4.3]). We use induction on the geometric intersection number of two simple closed curves $a$ and $b$ which bound a $2$-punctured disc on $S_{0,n}$. When they are disjoint there is nothing to prove. Let us then assume that $a$ and $b$ have geometric intersection $i(a,b)>0$. We claim that there is a simple closed curve $c$ (bounding a $2$-punctured disc) such that $i(a,c)=0$ and $i(c,b)<i(a,b)$.
Now, $c^{\prime}$ bounds a $k$-punctured disc $D$, where $k\geq 2$, which does not contain $a$. If $k=2$, then we take $c=c^{\prime}$ and we are done. Otherwise, we take for $c$ the boundary of a $2$-punctured disc $D^{\prime}$ contained in $D$ such that $c$ is the union of an arc contained in $c^{\prime}$ and an arc which is either part of $b\cap D$ or is disjoint from $b$. In both cases, we have that $i(c,b)\leq i(c^{\prime},b)<i(a,b)$ and, obviously, $i(a,c)=0$.
(ii): We proceed as above. Here, the curve $c$ is either nonseparating or it bounds a $2$-punctured disc but the proof above works without any essential change. ∎
2.7. Proof of Theorem 2.2 for $g(S)=0$
We proceed by induction on $n\geq 5$. The case $n=5$ was proved in Section 2.3 and serves as base for the induction. Let us then assume that Theorem 2.2 holds for $S_{0,n-1}$ and let us prove it for $S_{0,n}$.
Lemma 2.15.
Proof.
By [9, Theorem 6.7], for all hyperbolic surfaces $S$, there is a natural monomorphism:
which is induced by the identification of the vertices of ${\check{C}}_{P}(S)$ with the facets of ${\check{C}}(S)$.
Let us briefly recall the proof of this theorem. For $d(S)=0$, we have that ${\check{C}}_{P}(S)_{0}={\check{C}}(S)_{0}$ and $\dim{\check{C}}(S)=0$ and so the statement is obvious. For $d(S)>1$, from [9, Lemma 6.6, (ii)]), it follows that the continuous automorphisms of ${\check{C}}_{P}(S)$ preserve Farey subgraphs and so there is a continuous action of $\operatorname{Aut}({\check{C}}_{P}(S))$ on the profinite set of Farey subgraphs of ${\check{C}}_{P}(S)$. These are parameterized by the profinite set of (d(s)2)(d(s)-2)-simplices of ${\check{C}}(S)$, thus there is an induced continuous action on the latter set. By [9, Lemma 6.6, (iii)]), this continuous action induces a monomorphism $\operatorname{Aut}({\check{C}}_{P}(S))\hookrightarrow\operatorname{Aut}({\check{C}}^{\ast}(S))$, where ${\check{C}}^{\ast}(S)$ is the dual graph of ${\check{C}}(S)$ (cf.[9, Definition 3.9]). By [9, Lemma 6.5], we then have that $\operatorname{Aut}({\check{C}}^{\ast}(S))=\operatorname{Aut}({\check{C}}(S))$ and the conclusion follows.
Thus, after possibly composing the given automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$ with an element in the image of $\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))$, we can assume that its image $\check{\Theta}_{P}(\phi)$ in $\operatorname{Aut}({\check{C}}(S_{0,n}))$ preserves the 0-simplex $\{{\gamma}\}\in{\check{C}}(S_{0,n})$ and so $\phi$ preserves the subgraph $L_{\gamma}$ of ${\check{C}}_{P}(S_{0,n})$. Since, by Lemma 2.9, we have that $L_{\gamma}\cong{\check{C}}_{P}(S_{\gamma})$, from the induction hypothesis, it follows that, if the automorphism $\phi$ preserves the orientation of some profinite Farey subgraph of $L_{\gamma}$, then $\phi$ preserves the orientation of all profinite Farey subgraphs ${\widehat{\operatorname{F}}}_{\sigma}$ such that ${\gamma}\in{\sigma}$. ∎
By (i) of Lemma 2.14, the curve complex $C_{b}(S_{0,n})$ is connected. Thus, there is a set ${\gamma}_{1},\dots,{\gamma}_{k}$ of simple closed curves on $S_{0,n}$ bounding a $2$-punctured disc such that any two representatives $\sigma$} and $\sigma$$^{\prime}$ of the set of orbits $O(S_{0,n})$ are contained in a chain $\operatorname{Lk}({\gamma}_{1}),\dots,\operatorname{Lk}({\gamma}_{k})$ with the property that the intersection $L_{{\gamma}_{i}}\cap L_{{\gamma}_{i+1}}$, for $1\leq i\leq k-1$, contains at least a profinite Farey subgraph. Lemma 2.15 and a simple induction then imply that an automorphism $\phi\in\operatorname{Aut}({\check{C}}_{P}(S_{0,n}))$, which preserves the orientation of ${\widehat{\operatorname{F}}}_{\sigma}$, also preserves the orientation of ${\widehat{\operatorname{F}}}_{{\sigma}^{\prime}}$. This completes the proof of Theorem 2.2 for $g(S)=0$.
2.8. Proof of Theorem 2.2 for $g(S)\geq 1$
Let us first consider the case $g(S)=1$. Here, we need to use the curve complex $C_{0b}(S_{1,n})$ instead of the curve complex $C_{b}(S_{0,n})$ and (ii) instead of (i) of Lemma 2.14. We then proceed by induction on $n\geq 2$.
The base of the induction is provided by the case $S=S_{1,2}$ proved above. The induction step essentially proceeds as in the case $g(S)=0$ (cf. Section 2.7). The only difference is that, if ${\gamma}_{i}$, for $1\leq i\leq k$, is a nonseparating simple closed curve on $S$, then the link $\operatorname{Lk}({\gamma}_{i})$ is isomorphic to the procongruence curve complex of a genus 0 surface, so that, in this case, instead of the induction hypothesis, we need to use the genus 0 case of Theorem 2.2, which we already proved.
For $g(S)\geq 2$, we proceed by induction on the genus where the base of the induction is the genus $1$ case proved above. The relevant curve complex here is the complex of nonseparating curves $C_{0}(S)$, which, for $g(S)\geq 2$, (cf. [13, Theorem 4.4]) is connected. The rest of the argument proceeds as in the previous cases.
2.9. A rigidity criterion
From Theorem 2.2, it follows that, for $d(S)>1$, there is a natural isomorphism:
(8) $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S)).$
From this isomorphism, we will derive a characterization of those elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ which are induced by an inner automorphism of ${\check{\Gamma}}^{\pm}(S)$. Before we state the result, we need to make the following remark:
Remark 2.16.
In the group-theoretic realization of the procongruence curve complex ${\check{C}}(S)$ which we described in Remark 2.10, the vertices of the procongruence pants complex ${\check{C}}_{P}(S)$ are identified with the set $\{{\hat{\mathrm{I}}}_{\sigma}\}_{{\sigma}\in{\check{C}}(S)_{d(S)-1}}$ of inertia groups of $\mathrm{P}{\check{\Gamma}}(S)$ of maximal rank. The natural faithful continuous action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the curve complex ${\check{C}}(S)$ (cf. Lemma 2.11) then induces a continuous faithful action of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ on the vertex set ${\check{C}}_{P}(S)_{0}$ of the procongruence pants complex.
We have:
Theorem 2.17.
For the proof, we will need the following simple lemma in group theory:
Lemma 2.18.
Let $1\to H\to G\to L\to 1$ be a short exact sequence of groups and let $f$ be an automorphism of $H$ and suppose that:
@@ 2.9. A rigidity criterion
For an element $f\in\operatorname{Aut}(H)$, we denote by $\bar{f}$ its image in $\operatorname{Out}(H)$. Let then $\mathrm{Comp}(L,H)$ be the subgroup of $\operatorname{Aut}(L)\times\operatorname{Aut}(H)$ formed by the pairs $(\psi,f)$ such that, for all $\alpha\in L$, there holds (in $\operatorname{Out}(H)$):
@@
Since, by hypothesis (i), the center of $H$ is trivial, according to Wells’ exact sequence (cf. [40, Theorem]), there is a canonical isomorphism:
@@
Let $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ be an element satisfying the hypotheses of the theorem. Since, for $S=S_{0,5}$, there is only one topological type of both vertices and edges of the pants complex $C_{P}(S)$, the action of ${\check{\Gamma}}(S)$ on the set of oriented edges of ${\check{C}}_{P}(S)$ is transitive. Therefore, for any edge $e=\{\alpha_{0},\alpha_{1}\}$ of ${\check{C}}_{P}(S)$, there is an element $x_{e}\in{\check{\Gamma}}(S)$ such that $\alpha_{i}=x_{e}\cdot v_{i}\cdot x_{e}^{-1}$, for $i=0,1$ (cf. Remark 2.16).
Let us consider the short exact sequence $1\to\mathrm{P}{\check{\Gamma}}(S)\to{\check{\Gamma}}(S)\to\Sigma_{n}\to 1$, where we put $n:=n(S)$. As it is well known, ${\mathcal{M}}_{0,3}$ is a point and, for $n\geq 4$, the moduli space ${\mathcal{M}}_{0,n}$ is naturally isomorphic to the configuration space of n3n-3 ordered points on ${\mathbb{P}}^{1}\smallsetminus\{0,1,\infty\}$, since $\operatorname{Aut}({\mathbb{P}}^{1})$ acts simply transitively on triples of distinct points of ${\mathbb{P}}^{1}$. Hence, the group which in [18, Corollary C] is denoted by $\Pi_{n}$ is isomorphic to $\mathrm{P}{\check{\Gamma}}(S)$, so that [18, Corollary C] implies that the outer action of $\Sigma_{n}$ on $\mathrm{P}{\check{\Gamma}}(S)$, associated to the short exact sequence above, identifies $\Sigma_{n}$ with a normal subgroup of $\operatorname{Out}(\mathrm{P}{\check{\Gamma}}(S))$.
In particular, the automorphism $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ satisfies the second hypothesis of Lemma 2.18. Since $\mathrm{P}{\check{\Gamma}}(S)$ is center-free and the outer representation $\Sigma_{n}\to\operatorname{Out}(\mathrm{P}{\check{\Gamma}}(S))$ is faithful, the other two hypotheses are also satisfied. Therefore, by Lemma 2.18, $f$ extends to an automorphism $\tilde{f}$ of ${\check{\Gamma}}(S)$ which preserves procyclic inertia groups (cf. Definition 1.1) and so is contained in $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ (cf. Definition 1.2).
There is then a series of identities, for $i=0,1$:
$f(\alpha_{i})=\tilde{f}(\alpha_{i})=\tilde{f}(x_{e}\cdot v_{i}\cdot x_{e}^{-1})=\tilde{f}(x_{e})\cdot\tilde{f}(v_{i})\cdot\tilde{f}(x_{e})^{-1}=\tilde{f}(x_{e})\cdot f(v_{i})\cdot\tilde{f}(x_{e})^{-1},$
so that $\{f(\alpha_{0}),f(\alpha_{1})\}=\operatorname{inn}\tilde{f}(x_{e})(\{f(v_{0}),f(v_{1})\})\in{\check{C}}_{P}(S)_{1}$. Therefore, the continuous action of the automorphism $f$ on the profinite set of vertices of ${\check{C}}_{P}(S)$ extends to a continuous action on the procongruence pants complex ${\check{C}}_{P}(S)$. The conclusion then follows from the isomorphism (8).
@@
As for the proof of the case $S=S_{1,2}$ of Theorem 2.2 (cf. Section 2.4), the idea is to reduce to the genus 0 case. However, as in the proof of Theorem 2.2, this is complicated by the fact that, although the curve complexes $C(S_{1,2})$ and $C(S_{0,5})$ are naturally isomorphic, there is not even a map between the corresponding pants graphs. To deal with this case, we will then need first to prove a series of lemmas:
@@
With the notation of Section 2.5, for $\gamma$} a simple closed curve on $S_{1,2}$, by [9, Corollary 4.12], there is a natural isomorphism $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})\cong\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}$ and, by [9, Theorem 4.10], there are exact sequences:
$1\rightarrow\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{{\gamma}}\to\{\pm 1\}\hskip 14.22636pt\mbox{ and }\hskip 14.22636pt1\rightarrow\hat{\mathrm{I}}_{\gamma}\to\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}\to\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\to 1.$
@@
After possibly composing with an inner automorphism, we can assume that a given element $f\in\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ is such that ${\gamma}^{\prime}:=f({\gamma})$ also belongs to $C(S_{1,2})_{0}\subset{\check{C}}(S_{1,2})_{0}$. Since $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$ identifies with the subgroup of the centralizer $Z_{\mathrm{P}{\check{\Gamma}}(S_{1,2})}({\hat{\mathrm{I}}}_{\gamma})$ topologically generated by profinite Dehn twists, we have that $f(\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}})=\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}^{\prime}}$. By [9, Theorem 4.14], the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ is the center of $\mathrm{P}{\check{\Gamma}}(S_{1,2})_{\vec{{\gamma}}}$, so that $f$ induces an isomorphism $\bar{f}\colon\thinspace\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\stackrel{{\scriptstyle\sim}}{{\to}}\mathrm{P}{\check{\Gamma}}((S_{1,2})_{{\gamma}^{\prime}})$.
@@
For $\gamma$} separating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})\cong\widehat{\operatorname{SL}(2,\mathbb{Z})}$ while, for $\gamma$} nonseparating, we have that $\mathrm{P}{\check{\Gamma}}((S_{1,2})_{\gamma})$ is a free group in two generators. The latter profinite group is torsion-free while the former is not. Thus, $\gamma$$^{\prime}$ has the same topological type of $\gamma$}.
@@
This proves the first part of the lemma. Let us then observe that, by Lemma 2.6, $\mathrm{P}{\check{\Gamma}}(S_{0,5})$ identifies with the normal subgroup of $\mathrm{P}{\check{\Gamma}}(S_{1,2})$ topologically generated by squares of nonseparating Dehn twists. By the previous part of the proof, elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))$ preserve this subgroup and so there is a homomorphism as claimed in the lemma.
In order to prove that this homomorphism is injective, let us observe that the natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{1,2}))$ (cf. Lemma 2.11) factors through the given homomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{1,2}))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))$ and the natural monomorphism $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S_{0,5}))\hookrightarrow\operatorname{Aut}({\check{C}}(S_{0,5}))$, via the isomorphism ${\check{C}}(S_{1,2})\cong{\check{C}}(S_{0,5})$. This concludes the proof of the lemma. ∎
@@
Since all the groups involved are center-free, for $n\geq 4$, there is a series of natural isomorphisms:
@@
(9) $\operatorname{Inn}({\check{\Gamma}}(S_{0,n}))\left/\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S_{0,n}))\right.\cong{\check{\Gamma}}(S_{0,n})\left/\mathrm{P}{\check{\Gamma}}(S_{0,n})\right.\cong\Sigma_{n}.$